OpenStudy (anonymous):

A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 a. Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation.

OpenStudy (anonymous):

i really need help with this i have been working on it for about 2 hours

OpenStudy (anonymous):

This is a poorly phrased question. However, we can follow the instructions and answer it. I choose to multiply the first equation by 3/2, then add it to the second equation. We will take that sum and the first equation to be our new system of equations.

OpenStudy (anonymous):

Multiply both sides by 3/2\[2x+7y=1 \implies 3x+\frac{21y}{2}=\frac{3}{2}\]Add the equations together\[3x-3x+\frac{21y}{2}-4y=\frac{3}{2}+5 \implies \frac{13y}{2}=\frac{13}{2} \implies y=1\]So our system of equations is\[2x+7y=1\]\[y=1\]

OpenStudy (anonymous):

From the above, the only step left to solve the system is to substitute the second equation (y = 1) into the first equation and solve. If we substitute, we get\[2x+7=1 \implies 2x=-6 \implies x=-3\]So the solution of the system is\[(x,y)=(-3,1)\]